OpenStudy (habbababba1):

will fan and medal!!!....solve this system of linear equations. Seperate the x- and y- values with a comma

OpenStudy (habbababba1):

-10x=-88-3y 6x=24-3y

OpenStudy (habbababba1):

@Vocaloid

OpenStudy (habbababba1):

@rebeccaxhawaii

OpenStudy (habbababba1):

@inkyvoyd

OpenStudy (salazarblack):

this is easy ..just multiply one equation by -1

OpenStudy (salazarblack):

and then -3y and -3y subtracts

OpenStudy (salazarblack):

and afterwards you sum the two equations and find x

OpenStudy (salazarblack):

when you do that you put the value of x in some of the 2 equations and get the value for y

OpenStudy (habbababba1):

do -1for x?

OpenStudy (salazarblack):

-10x=-88-3y 6x=24-3y / (-1)

OpenStudy (salazarblack):

-10x=-88-3y -6x=-24+3y

OpenStudy (salazarblack):

now you sum the two and get -10x-6x=-88-24

OpenStudy (salazarblack):

from this you can get the x value